OpenStudy (anonymous):

Question : 5 Signals : \[\large \color{green}{y(t) = \frac{1}{T} \int\limits_{t - \frac{T}{2}}^{t + \frac{T}{2}}x(t) \cdot dt}\] \(\text{How this system is Time-Invariant ??}\)

OpenStudy (anonymous):

Am I doing it right??

OpenStudy (anonymous):

\[x_1(t) \longrightarrow x(t - t_0)\] \[y_1(t) \longrightarrow \frac{1}{T} \int\limits_{t - \frac{T}{2}}^{t+ \frac{T}{2}} x_1(t) dt \implies \frac{1}{T} \int\limits_{t - \frac{T}{2}}^{t+ \frac{T}{2}} x(t - t_0) dt\]

OpenStudy (anonymous):

Put : \(t - t_0 = \lambda\) such that \(dt = d \lambda\)

OpenStudy (anonymous):

\[y_1(t) \longrightarrow \frac{1}{T} \int\limits_{t - t_0- \frac{T}{2}}^{t - t_0+ \frac{T}{2}} x(\lambda) d \lambda\]

OpenStudy (anonymous):

While : \[y(t - t_0) = \frac{1}{T} \int\limits_{t - t_0 -\frac{T}{2}}^{t - t_0+ \frac{T}{2}} x(t - t_0) dt\]

OpenStudy (anonymous):

So: \[y_1(t) \longrightarrow y(t-t_0)\] So, the system is Time-Invariant..

OpenStudy (anonymous):

To what extent I am right in doing these steps???

myininaya (myininaya):

well you use the sub \[t-t_0=\lambda\] So That would mean\[t=\lambda+t_0\]

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

Where you are saying??

OpenStudy (anonymous):

I mean where you are saying to replace it with \(\lambda + t_0\) ??

myininaya (myininaya):

Oh I guess you are saying that x(t) is equivalent to x(t-t_0)

myininaya (myininaya):

I'm just looking at your integral transformation from something in terms of t to something in terms of lambda

myininaya (myininaya):

why is t replaced with t-t_0?

myininaya (myininaya):

and not lambda+t_0

OpenStudy (anonymous):

Yes, \(x_1(t)\) = \(x(t - t_0)\)

myininaya (myininaya):

But the limits...why is t replaced with t-t_0

OpenStudy (anonymous):

It is for substitution I did..

myininaya (myininaya):

So t=t-t_0 is your substitution?

OpenStudy (anonymous):

\[t - t_0 = \lambda\] Now change limits..

myininaya (myininaya):

I'm just trying to figure out why the t was replaced with t-t_0

OpenStudy (anonymous):

No.. \(t - t_0 = \lambda\) is my substitution..

OpenStudy (anonymous):

Yes, it is for checking time invariance or variance of a system..

OpenStudy (anonymous):

1. First you check if your input was \(x(t - t_0)\), then your output is changing correspondingly with same amount or not..

OpenStudy (anonymous):

\[If \quad x(t ) \longrightarrow x(t -t_0) \quad then \quad y(t) \longrightarrow y(t-t_0)\]

OpenStudy (anonymous):

If output is changing by same amount as the input is, then the system is time invariant, and input and output relationship is maintained.. :)

OpenStudy (anonymous):

Are you getting Ma'm what I am trying to do here?? You can very well tell me if I am wrong anywhere according to you.. I am seriously not sure, how to check Time-Invariance of any system.. :(

myininaya (myininaya):

I'm looking to see if I can find a section on this in my calculus book.

OpenStudy (anonymous):

Oh wait.. It is from Signals And Systems..

OpenStudy (anonymous):

I am not sure if you will find it in your Calculus book or not.. :)

myininaya (myininaya):

Ok then I will google signals and systems integrals

OpenStudy (anonymous):

That will be nice.. :)

myininaya (myininaya):

Hey so that capital T outside is just a constant...It is not t right?

OpenStudy (anonymous):

Getting something to work on Ma'm?? I need to go now, time is not permitting me to stay here more.. :(

OpenStudy (anonymous):

Oh yes, it is just constant.. Capital T is everywhere constant..

OpenStudy (anonymous):

A signal is represented as : \(x(t)\), where t is time parameter.. As it is One Dimensional Signal, only t is variable here and others are constants..

myininaya (myininaya):

http://en.wikibooks.org/wiki/Signals_and_Systems/Time_Domain_Analysis We want to check to see if these two integrals are equal: \[\frac{1}{T}\int\limits_{t-\pi/2}^{t+\pi/2}x(t+d) dt \] \[\frac{1}{T}\int\limits_{t+d-\pi/2}^{t+d+\pi/2}x(t+d) dt \] and yes I used d instead to t_0 That is what I have come to so far... And we would need to use transformations to see if these are indeed equal.

OpenStudy (anonymous):

I don't think they are equal.. :(

myininaya (myininaya):

Well I would have to do transformations to see... But I think you already have gotten these integrals above.

OpenStudy (anonymous):

No, I have got something different..

myininaya (myininaya):

I don't see any difference.

myininaya (myininaya):

I have d instead of -t_0 this is just a constant

myininaya (myininaya):

anyways I do not think the first integral equals the second integral

OpenStudy (anonymous):

I mean: Look at my \(y_1(t)\) and \(y(t-t_0)\) expressions..

OpenStudy (anonymous):

Yes, that is what I think, Oh so, we think the same.. :P

OpenStudy (anonymous):

There \(y_1(t)\) and \(y(t-t_0)\) are equal I think..

myininaya (myininaya):

yes

myininaya (myininaya):

The only difference is instead of them using -t_0 they put +d this is not really a difference though

myininaya (myininaya):

wait wait I think I'm wrong y_1 =y_2 i think

myininaya (myininaya):

I will show you

OpenStudy (anonymous):

Difference? Which two expressions you are looking difference in??

myininaya (myininaya):

So we are comparing \[y_2=\frac{1}{T}\int\limits_{t-t_0-\pi/2}^{t-t_0+\pi/2} x(t-t_0) dt \\ \text{ and } y_1=\frac{1}{T}\int\limits_{t-\pi/2}^{t+\pi/2} x(t-t_0) dt \] I will go back and use -t_0 instead +d So you see which of your two integrals I'm comparing I will use the sub on the first integral \[\lambda=t-t_0 => d \lambda= d t \] so we have \[y_2(\lambda)=\frac{1}{T}\int\limits_{\lambda-\pi/2}^{\lambda+\pi/2}x(\lambda) d \lambda \] So if our input was t we would have \[y_2(t)=\frac{1}{T}\int\limits_{t-\pi/2}^{t+\pi/2}x(t) dt \] But you told me x(t)=x(t-t0) so y_1=y_2

myininaya (myininaya):

so they are time in-variant

OpenStudy (anonymous):

Wait, let me take me some time for viewing it properly.. :)

OpenStudy (anonymous):

That's the thing I was looking for.. :) Real help.. :) Thank you Ma'm.. :)

myininaya (myininaya):

I'm glad I could help. I guess it makes perfect sense?

OpenStudy (anonymous):

Just perfect, nothing greater than that.. :)

myininaya (myininaya):

nice

OpenStudy (anonymous):

I have many more with having doubts in to work on.. :) May be next time I will pick them.. :)